### v2

@@ 2 $f(R)$ Gravity
These modifications are crucial for testing $f(R)$ gravity with detectors like LIGO and Virgo, as they provide a means to distinguish between different gravity models based on the behavior of the observed GW signals. For instance, the absence of scalar polarization modes in the detected signals from binary black hole mergers has already placed stringent constraints on $f(R)$ gravity models [17]. Future observations with next-generation detectors, such as the Einstein Telescope and Cosmic Explorer, will further improve these constraints by providing higher sensitivity to the polarization and dispersion properties of gravitational waves. Moreover, the detection of the stochastic gravitational wave background (SGWB) could offer additional insights into the scalar mode, as its energy density and frequency spectrum are sensitive to the modifications introduced by $f(R)$ gravity [33].
@@ Acknowledgements
I would like to express my thanks for the advice and help I received from Andrew L. Miller, and for the very useful conversations with Şengül Kuru, Banu Şahin, Giovanni Montani, and Anne-Christine Davis. I would also like to acknowledge the support from the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology), and for their encouragement for young researchers.

### v3

@@ 2 $f(R)$ Gravity
These modifications are crucial for testing $f(R)$ gravity with detectors like LIGO and Virgo, as they provide a means to distinguish between different gravity models based on the behavior of the observed GW signals. For instance, the absence of scalar polarization modes in the detected signals from binary black hole mergers has already placed stringent constraints on $f(R)$ gravity models [17]. Future observations with next-generation gravitational wave detectors, such as the Einstein Telescope and Cosmic Explorer, will significantly enhance our ability to test modifications to general relativity. These advanced instruments will offer unprecedented sensitivity to key properties of gravitational waves, including their propagation speed and polarization content. In particular, multi-messenger observations—such as the coincident detection of gravitational waves and their electromagnetic counterparts—enable precise constraints on the propagation speed of gravitational waves, a test originally proposed by [34, 33]. In the context of general modified gravity theories, such as certain formulations of Horndeski gravity, the tensor modes of gravitational waves may propagate at speeds different from that of light. Current observations have already placed stringent limits on such deviations, as demonstrated in Ref. [35], which specifically constrains the propagation speed of tensor modes in Horndeski theory—a framework that includes $f(R)$ gravity as a special case.
In $f(R)$ gravity, both the tensor and scalar modes propagate at the speed of light, consistent with general relativity and with more general modified gravity theories that feature non-minimal coupling between curvature and matter [36]. Consequently, deviations in propagation speed do not serve as a distinguishing feature between $f(R)$ gravity and general relativity. Instead, the hallmark of $f(R)$ gravity lies in the presence of an additional scalar degree of freedom, which manifests as an extra polarization mode in gravitational waves, beyond the standard plus and cross polarizations of general relativity. This scalar mode, often referred to as a transverse breathing mode, can be probed through detailed polarization analysis of gravitational wave signals. Recent studies utilizing data from the LIGO and Virgo observatories have begun to constrain the amplitude of such scalar modes, providing observational insights into their potential presence [37, 38].
However, recent advancements have refined our understanding of the linear degrees of freedom in $f(R)$ gravity, particularly concerning the propagation of gravitational waves in maximally-symmetric space-times, such as flat Minkowski space. Initially, it was assumed that all $f(R)$ models propagate three polarization modes—two tensor modes associated with the graviton and one scalar mode linked to the scalaron. Contrarily, a significant class of $f(R)$ models, newly classified as “degenerate models”, has been shown to possess an empty linear spectrum in such backgrounds, meaning they do not propagate any gravitational waves at all [39, 40]. This class includes notable examples such as $f(R)=R^{2}$ and various power-law forms frequently utilized in cosmological applications. These findings [39, 40] overturn the earlier belief of universal propagation of three polarizations across all $f(R)$ theories, highlighting that the presence of gravitational wave modes is contingent upon the specific functional form of $f(R)$. Moreover, these degenerate models exhibit stability issues and other unphysical properties, casting doubt on their physical viability as consistent gravitational theories [39, 40]. This revelation underscores the necessity of carefully delineating between degenerate and non-degenerate $f(R)$ models when interpreting gravitational wave observations, as only the latter support the additional scalar polarization mode previously discussed.
@@ Acknowledgements
I would like to express my thanks for the advice and help I received from Andrew L. Miller, and for the very useful conversations with Şengül Kuru, Banu Şahin, Giovanni Montani, and Anne-Christine Davis. I would also like to acknowledge the support from the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology), and for their encouragement for young researchers. Additionally, I would also like to extend my gratitude to vaerious anonymous reviewers for their insightful comments and constructive feedback, which greatly contributed to improving the quality of this paper.